what is the value of X in the problem
x+2x+20+3x-40+4x-20=360

Answer by london maths tutor(243)

x+2x+20+3x-40+4x-20=360
(1+2+3+4)x+20-40-20 = 360
10x - 40 = 360
10x = 360 + 40
10x = 400
x = 400/10
x = 40